OpenStudy (lovelycharm):

Eric has been donating books he no longer needs to his local library. After 3 months of donating, he had 157 books left in his room. After 8 months of donating, he had 102 books left in his room. How many books per month did Eric donate on average between months 3 and 8?

OpenStudy (ddcamp):

In general: \[Average=\frac{ Total }{ Time }\] How many books did he donate in total? How much time passed?

OpenStudy (anonymous):

Total books = 157+3x Total books = 102+8x So: 157+3x=102+8x\[\rightarrow 5x=55 \rightarrow x=11\]

OpenStudy (anonymous):

Each month 11, so between 3 and 8 months (I supoose not included, 4 months) 4x11=44

OpenStudy (lovelycharm):

thank you